OpenStudy (anonymous):

What is this problem? Identity? No Solution? 1/7(14-7p)-2=-2(1/2p-3)+6

OpenStudy (anonymous):

Help?

OpenStudy (anonymous):

Does this make sense?

OpenStudy (anonymous):

If you need me to break it down some more I can certainly do that for you.

OpenStudy (anonymous):

Wouldn't the problem end up like this 2-p+2=-p-6+6?

OpenStudy (anonymous):

You need to watch your signs. Remember, a negative times a negative is a positive. Also, a positive times a negative is a negative. Review your process. And yes, you're correct, the numbers can be set up that way too. Though you will still get the same issue of a number equaling another number. I've detailed it a bit more for you.

jhonyy9 (jhonyy9):

on the left side will remain 2-p-2 =p yes ?

jhonyy9 (jhonyy9):

on the right side -p+3+6=9-p so p=9-p 2p=9 p=9/2